Please could someone kindly help me???
This is my question:
The perimeter of a rectangle is 140in. The length is 28in. greater than the width. Find the width of the rectangle.
This is how I'm doing it could you tell me if its right???
2(x+28)=140
2x+56=140
2x=140-56
2x=84
2 2
x=42
42 Divide 2
=21 width
Please help:)
Thank you

Answer by checkley77(12844)

L=W+28
2L+2W=140
2(W+28)+2W=140
2W+56+2W=140
4W=140-56
4W=84
W=84/4
W=21 ANS. FOR THE WIDTH.
L=21+28=49 ANS. FOR THE LENGTH.
PROOF:
2*49+2*21=140
98+42=140
140=140